# Automaton Constrained Q-Learning

### v1

@@ 4.1 Augmented Product CMDP Formulation
Although the sum-of-costs formulation in (1) conveniently leverages the standard Bellman equation, it presents practical challenges in learning. Specifically, predicting the cumulative sum of future costs requires accurately modeling long-term dependencies, which increases variance and slows down convergence. This can be appropriate for maximizing reward but unnecessarily complicated for safety. When dealing with safety. and in particular hard-safety [43], we are interesting in knowing whether the generated trajectory becomes unsafe at least once. Hence, this can be reformulated as a classification task. This approach bypasses the need for accurate regression by directly learning the decision boundary for state-action pairs leading to safety violations. To achieve this, we define a constraint function $c^{A}$ such that it takes negative values if a safety constraint is violated and positive values otherwise, and we apply a constraint based on the minimum over all future values of $c^{A}(s^{A}_{t},a_{t})$. Then, we simply learn to classify this minimum value as being either positive or negative. This formulation allows us to determine safety violations in a single step, avoiding the complexity of cumulative cost prediction. A major advantage of this formulation is that it removes the need for manual tuning of the violation limit $L\in\mathbb{R}$, which depends on the range of possible costs. Instead, by directly classifying actions as leading to a safety violation, we reduce the range of $L$ to [1,1][-1,1]. This normalization ensures robustness across tasks and allows for a simple, task-agnostic choice of $L=0$. Directly defining $c^{A}(\langle s,g^{+},q\rangle,a)=\rho(s,S(q))$ using the robustness function for $S(q)$ (see Section 3) satisfies this formulation and corresponds to the safety constraints for any STL-specified task.
Now, we can define our new objective as finding the optimal stationary policy $\pi^{*}:\mathcal{S}\to P(\mathcal{A})$ that maximizes the expected total discounted return, as defined in (1), while keeping the expected minimum safety above the safety limit $L$:
$\mathop{\mathbb{E}}_{\tau\sim\pi}\left[\min_{t\in[0,\infty]}c^{A}(s_{t},a_{t})\right]>L,$ (2)
@@ 4.2 Automaton Constrained Q-Learning (ACQL)
In what follows, we provide an overview of the ACQL algorithm for learning
$\pi^{*}(s^{A})=\operatorname*{arg\,max}_{a\;:\;Q^{c*}(s^{A},a)>L}Q^{r*}(s^{A},a),$
where $Q^{r*}=Q^{r}_{\pi^{*}}$ is the optimal state-action value function and $Q^{c*}=Q^{c}_{\pi^{*}}$ is the optimal state-action safety function. In order to learn these optimal functions, we define the models $Q^{r}_{\psi}:\mathcal{S}^{A}\times\mathcal{A}\rightarrow\mathbb{R}$ and $Q^{c}_{\theta}:\mathcal{S}^{A}\times\mathcal{A}\rightarrow[-1,1]$ parameterized by $\psi$ and $\theta$. While exploring the environment via an epsilon-greedy policy, $Q^{r}_{\psi}$ is updated to minimize the loss
$L^{r}_{j,i}(\psi)=\mathop{\mathbb{E}}_{(s^{A}_{t},a_{t})\sim\mathcal{B}_{i}}\left[\left(y^{r}_{t}-Q^{r}_{\psi}(s^{A}_{t},a_{t})\right)^{2}\right],\ \ \text{with}\ \ y^{r}_{t}=r_{t}+\gamma Q^{r}_{\bar{\psi}}(s^{A}_{t+k},\pi_{j}(s^{A}_{t+k})).$ (3)
The inclusion of subgoals in the states $s^{A}_{t}\in\mathcal{B}_{i}$ allows us to further accelerate learning using techniques such as HER [2], which allows $Q^{r}_{\theta}$ to improve from failed attempts at progressing through the task. Similarly, $Q^{c}_{\theta}$ is trained with the target
@@
which is derived from the Bellman principle of optimality for expected minimum cost objectives given in [16]. The targets, $y^{r}$ and $y^{c}$, are computed using target parameters, $\bar{\psi}$ and $\bar{\theta}$, which are updated towards their corresponding main parameters each step with an interpolation factor $\lambda$. Crucially, we schedule the discount factor $\gamma_{c}$ throughout training to asymptotically approach $1.0$, which is necessary for convergence to $Q^{c*}$.
@@
The step sizes $a(n)$, $b(n)$, and $c(n)$ create a stochastic approximation algorithm on three timescales [44]. Based on the contraction property of (4), the fastest updating $Q^{c}_{n}$ asymptotically tracks the correct safety function for the policy determined by $Q^{r}_{n}$ and $\gamma_{c_{n}}$. On the slower timescale ensured by the definition of $b(n)$ and $a(n)$, $Q^{r}_{n}$ converges to the fixed point as determined by the relatively static $\gamma_{c_{n}}$, due to the contraction property of (3). Finally, as $\gamma_{c}$ converges on the slowest timescale, the fixed points for the two faster timescales approach $Q^{r*}$ and $Q^{c*}$ at $\gamma_{c}=1$. ∎
In Proposition 1, the notation $b(n)\in o(a(n))$ means that $b(n)/a(n)\to 0$ asymptotically. The assumption of a finite MDP is standard in the literature, but relaxing this assumption is possible by applying fixed point theorems for infinite dimensional spaces [45]. After convergence, the final policy will behave optimally for the reward $r$ and (if possible) never incur any cost, thereby respecting the LTL safety constraint. A complete description of our algorithm, model implementation details, and a full proof of convergence, with reference to similar arguments used in prior work [46, 47, 34], are provided in the appendix.
@@ 5 Experiments
In this section, we justify the design of ACQL and evaluate its effectiveness. We conducted a comparative analysis of ACQL against established baselines in RL from LTL specifications, including RMs [12] and the LOF [13]. We demonstrate our method’s real-world applicability by solving more complex LTL tasks with a 6-DOF robot arm in a storage cabinet environment. Lastly, we performed an ablation study to clarify how our subgoal-including product CMDP, in combination with HER [2], and our minimum-safety constraint formulation contribute to ACQL’s performance.
@@ 5.1 Comparative Analysis
In selecting our baselines, we chose to not compare against standard Safe RL methods or offline policy learning methods. Safe RL methods typically address stationary safety constraints, and do not generalize to the non-stationary safety requirements expressible in LTL specifications. Likewise, offline methods assume access to a fixed dataset and are not designed for the online learning setting we address. Instead, we compared our algorithm against two methods for online RL from LTL specifications: Counterfactual Experiences for Reward Machines with Reward Shaping (CRM-RS) [30] and the LOF [13]. CRM-RS employs an RM to define the task and dispense shaped rewards that incentivize task completion. RM states in non-accepting sink components are treated as terminal states, implementing a rollout-terminating strategy to enforce safety constraints. The LOF is a framework for learning hierarchical policies to satisfy LTL tasks. It learns a separate policy, referred to as a logical option, for satisfying every subgoal-proposition in the task. When training each logical option, each step that violates the task’s safety propositions receives a large reward penalty (set to $R_{s}=-1000$ based on their implementation) to discourage unsafe behavior. These policies are then composed with a higher-level policy obtained using value iteration over a discrete state space derived from the task automaton and all the subgoal states of the environment. Both these approaches can handle general LTL specifications and, to our knowledge, represent the most relevant baselines to compare our algorithm against. Furthermore, we note that RMs and the LOF cannot be easily enhanced with high-performing GCRL and Safe RL techniques without adopting substantial modifications to their learning algorithms and product MDP formulation. Our method effectively integrates these components in a single framework that enables learning from general logical objectives in complex settings, as demonstrated in the following tasks.
@@
We chose five distinct LTL tasks and three different agents to facilitate a thorough comparison between our algorithm and the two baseline methods. All environment simulation was done within the Brax physics simulator [48]. The agents used in our experiments are a 2D PointMass, a Quadcopter, and an 8-DOF Ant quadruped. Additional details for these environments are given in the appendix. We used open environments without physical obstacles for our experiments. Instead, obstacles were introduced solely through the task specification. This was done to demonstrate the effectiveness of our algorithm in avoiding obstacles based purely on feedback from the task specifications, without confounding effects such as physically restricted motion. The LTL tasks used in the evaluation include: (1) a two-subgoal sequential navigation task between opposite corners of the environment ($\lozenge(g_{1}\wedge\circ(\lozenge g_{2}))$), (2) a two-subgoal branching navigation task between opposite corners of the environment ($\lozenge g_{1}\wedge\lozenge g_{2}$), (3) a single-goal navigation task constrained by an unsafe region ($\lozenge g_{1}\wedge\square\neg o_{1}$), (4) a two-subgoal navigation task with a disappearing safety constraint ($\neg o_{1}\mathcal{U}g_{1}\wedge\circ\lozenge g_{2}$), and (5) an infinitely-looping navigation task with a persistent safety constraint ($\square\lozenge(g_{1}\wedge\circ\lozenge g_{2})\wedge\square\neg o_{1}$). The starting position, obstacle and subgoal configurations for each environment are shown in Figure LABEL:fig:environments.
@@ Results
The results for our three simulated environments and five task types are shown in Table LABEL:tab:task-experiments. All results are reported for the final policy obtained after 5 million environment interactions with five different seeds. For LOF, each individual logical option was afforded its own 5 million training steps for fairness. For each seed, we evaluated the final policy in 16 randomly initialized episodes lasting 1000 steps. We report the reward, which corresponds to the number of steps spent near the final goal of the task automaton, and the success rate ($S.R.$), which is the proportion of policy rollouts that completely satisfied the task specification for all 1000 steps. Because the robustness of rollouts for the final “Loop” task cannot be meaningfully evaluated over a finite trajectory [49], we instead compute the success rate by examining the proportion of rollouts that are never unsafe and successfully complete at least 1 full loop within the episode.
@@ 5.3 Ablative Analysis
Our second ablation demonstrates the benefit of our minimum safety constraint (2) over the discounted sum-of-cost framework (1) for accurately learning the state-action safety function. We modified our CMDP to provide positive costs when there was a safety violation, trained $Q^{c}_{\theta}$ to predict discounted sums of costs, and constrained $\pi$ to choose actions below an upper limit on total cost $L$. This limit was chosen based on the best performing value $L\in\{0,10,40,60\}$. We report the difference in performance over only the tasks that involved safety constraints and observe that our CMDP formulation is critical to the performance of our algorithm. We also show a visualization of the policy differences that result from this ablation in Figure 3. We observe that in our formulation (left), ACQL can more effectively learn the safety constraint, which substantially improves the quality of $Q^{r}_{\psi}$ (bottom). Note that the differing scale for $Q^{c}_{\theta}$ in our formulation (top left) results in a yellow safe region, while the standard formulation depicts low-cost, safe regions in blue.

### v2

@@ 4.1 Augmented Product CMDP Formulation
Although the sum-of-costs formulation in (1) conveniently leverages the standard Bellman equation, it presents practical challenges in learning. Specifically, predicting the cumulative sum of future costs requires accurately modeling long-term dependencies, which increases variance and slows down convergence. This can be appropriate for maximizing reward but unnecessarily complicated for safety. When dealing with safety. and in particular hard-safety [wang2023hardconstraints], we are interesting in knowing whether the generated trajectory becomes unsafe at least once. Hence, this can be reformulated as a classification task. This approach bypasses the need for accurate regression by directly learning the decision boundary for state-action pairs leading to safety violations. To achieve this, we define a constraint function $c^{A}$ such that it takes negative values if a safety constraint is violated and positive values otherwise, and we apply a constraint based on the minimum over all future values of $c^{A}(s^{A}_{t},a_{t})$. Then, we simply learn to classify this minimum value as being either positive or negative. This formulation allows us to determine safety violations in a single step, avoiding the complexity of cumulative cost prediction. A major advantage of this formulation is that it removes the need for manual tuning of the violation limit $\mathcal{L}\in\mathbb{R}$, which depends on the range of possible costs. Instead, by directly classifying actions as leading to a safety violation, we reduce the range of $\mathcal{L}$ to [1,1][-1,1]. This normalization ensures robustness across tasks and allows for a simple, task-agnostic choice of $\mathcal{L}=0$. Directly defining $c^{A}(\langle s,g^{+},q\rangle,a)=\rho(s,S(q))$ using the robustness function for $S(q)$ (see Section 3) satisfies this formulation and corresponds to the safety constraints for any STL-specified task.
Now, we can define our new objective as finding the optimal stationary policy $\pi^{*}:\mathcal{S}\to P(\mathcal{A})$ that maximizes the expected total discounted return, as defined in (1), while keeping the expected minimum safety above the safety limit $\mathcal{L}$:
$\mathop{\mathbb{E}}_{\tau\sim\pi}\left[\min_{t\in[0,\infty]}c^{A}(s_{t},a_{t})\right]>\mathcal{L},$ (2)
@@ 4.2 Automaton Constrained Q-Learning (ACQL)
Algorithm 1 Automaton Constrained Q-Learning
1:An MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},d_{0},r,\gamma)$, an STL specification $\phi\in\Phi$, a safety limit $\mathcal{L}\in[-1,1]$, a learning rate $\alpha$, an interpolation factor $\lambda$
[…]
8:for $j=1,\ldots,N$ do
9:  $\gamma_{c}\leftarrow\textsc{SafetyGammaScheduler}(j)$
10:  $\pi_{j}(s^{A})\leftarrow\operatorname*{arg\,max}_{a:Q^{c}_{\theta}(s^{A},a)>\mathcal{L}}Q^{r}_{\psi}(s^{A},a)$
11:  $\tau\leftarrow\textsc{GetTrajectory}(\mathcal{M}^{A},\pi_{j})$
12:  $R\leftarrow R\cup\tau$.
13:  $\mathcal{B}^{\tau}_{j}\sim R$
14:  $\mathcal{B}^{\tau}_{j}\leftarrow\textsc{Relabel}(\mathcal{B}^{\tau}_{j})$
15:  for $i=1,\ldots,M$ do
16:   $\mathcal{B}_{i}\leftarrow\{(s^{A}_{t},a_{t},r^{A}_{t},c^{A}_{t},s^{A}_{t+1})\}\sim\mathcal{B}^{\tau}_{j}$
17:   $\theta\leftarrow\theta-\alpha\nabla_{\theta}L^{c}_{j,i}(\theta)$
18:   $\psi\leftarrow\psi-\alpha\nabla_{\psi}L^{r}_{j,i}(\psi)$
19:   $\bar{\theta}\leftarrow(1-\lambda)\bar{\theta}+\lambda\theta,\bar{\psi}\leftarrow(1-\lambda)\bar{\psi}+\lambda\psi$
20:  end for
21:end for
@@
In what follows, we provide an overview of the ACQL algorithm, whose pseudocode is shown in Algorithm 1. First, the input STL specification $\phi$ is translated into an automaton $A$ (Line 1) that is partitioned to produce the mappings $S$ and $G$ (Line 2). These mappings and the automaton are combined with the input MDP $\mathcal{M}$ to create the augmented CMDP in Line 3. Using the augmented CMDP, ACQL can proceed to learn the optimal policy
$\pi^{*}(s^{A})=\operatorname*{arg\,max}_{a\;:\;Q^{c*}(s^{A},a)>\mathcal{L}}Q^{r*}(s^{A},a),$
by learning the optimal state-action value function $Q^{r*}=Q^{r}_{\pi^{*}}$ and the optimal state-action safety function $Q^{c*}=Q^{c}_{\pi^{*}}$. In order to learn these optimal functions, we define the models $Q^{r}_{\psi}:\mathcal{S}^{A}\times\mathcal{A}\rightarrow\mathbb{R}$ and $Q^{c}_{\theta}:\mathcal{S}^{A}\times\mathcal{A}\rightarrow[-1,1]$ parameterized by $\psi$ and $\theta$ (Line 4). Their initial parameters are copied to initialize the target parameters $\bar{\theta}$ and $\bar{\psi}$ in Line 5, and an empty replay buffer is initialized in Line 6.
The algorithm proceeds to iterate for $N$ epochs indexed by $j$. At each epoch, we obtain a value for the safety discount factor $\gamma_{c}$ in Equation (4), which asymptotically approaches $1.0$ as training progresses (Line 8). The policy $\pi_{j}$ for the epoch is defined to select the most-rewarding action according to $Q^{r}_{\psi}$ constrained by $Q^{c}_{\theta}$ (Line 9). A trajectory $\tau$ is collected according to an epsilon-greedy version of this policy (Line 10), and this trajectory is added to the replay buffer (Line 11). From this replay buffer, a batch of trajectories $\mathcal{B}^{\tau}_{j}$ is sampled every epoch (Line 12). The inclusion of subgoals in the states $s^{A}_{t}$ of the sampled mini-batch $\mathcal{B}_{i}$ allows us to further accelerate learning using relabeling techniques such as HER [andrychowicz2017her], which allows $Q^{r}_{\theta}$ to improve from failed attempts at progressing through the task. Mini-batches of transitions $\mathcal{B}_{i}$ are subsequently sampled from the relabeled trajectories (Line 15) and used to update $Q^{r}_{\psi}$ for minimizing the loss
$L^{r}_{j,i}(\psi)=\mathop{\mathbb{E}}_{(s^{A}_{t},a_{t})\sim\mathcal{B}_{i}}\left[\left(y^{r}_{t}-Q^{r}_{\psi}(s^{A}_{t},a_{t})\right)^{2}\right],\ \ \text{with}\ \ y^{r}_{t}=r^{A}_{t}+\gamma Q^{r}_{\bar{\psi}}(s^{A}_{t+1},\pi_{j}(s^{A}_{t+1})).$ (3)
Similarly, $Q^{c}_{\theta}$ is trained with the target
@@
which is derived from the Bellman principle of optimality for expected minimum cost objectives given in [fisac2019bridging]. The targets, $y^{r}$ and $y^{c}$, are computed using target parameters, $\bar{\psi}$ and $\bar{\theta}$, which are updated towards their corresponding main parameters each step with an interpolation factor $\lambda$ (Line 18). Crucially, we schedule the discount factor $\gamma_{c}$ throughout training to asymptotically approach $1.0$, which is necessary for convergence to $Q^{c*}$. Additional details for the subroutines used in Algorithm 1 are provided in the appendix.
@@
The step sizes $a(n)$, $b(n)$, and $c(n)$ create a stochastic approximation algorithm on three timescales [borkar2008stochasticapproximation]. Based on the contraction property of (4), the fastest updating $Q^{c}_{n}$ asymptotically tracks the correct safety function for the policy determined by $Q^{r}_{n}$ and $\gamma_{c_{n}}$. On the slower timescale ensured by the definition of $b(n)$ and $a(n)$, $Q^{r}_{n}$ converges to the fixed point as determined by the relatively static $\gamma_{c_{n}}$, due to the contraction property of (3). Finally, as $\gamma_{c}$ converges on the slowest timescale, the fixed points for the two faster timescales approach $Q^{r*}$ and $Q^{c*}$ at $\gamma_{c}=1$. ∎
In Proposition 1, the notation $b(n)\in o(a(n))$ means that $b(n)/a(n)\to 0$ asymptotically. The assumption of a finite MDP is standard in the literature, but relaxing this assumption is possible by applying fixed point theorems for infinite dimensional spaces [granas2003fixed]. After convergence, the final policy will behave optimally for the reward $r$ and (if possible) never incur any cost, thereby respecting the LTL safety constraint. A complete description of our algorithm, model implementation details, and a full proof of convergence, with reference to similar arguments used in prior work [ma2022jointsynthesissafetycertificate, chow2017riskconstrainedreinforcementlearningpercentile, yu2022reachabilityconstrainedreinforcementlearning], are provided in the appendix.
@@ 5 Experiments
In this section, we justify the design of ACQL and evaluate its effectiveness. We conducted a comparative analysis of ACQL against established baselines in RL from LTL specifications, including RMs [toroicarte2018RewardMachines] and the LOF [araki2021logicaloptionsframework]. We demonstrate our method’s real-world applicability by solving more complex LTL tasks with a 6-DOF robot arm in a storage cabinet environment. Lastly, we performed an ablation study to clarify how our subgoal-including product CMDP, in combination with HER [andrychowicz2017her], and our minimum-safety constraint formulation contribute to ACQL’s performance.
@@ 5.1 Comparative Analysis
In selecting our baselines, we chose to not compare against standard Safe RL methods, offline policy learning methods, and multi-task LTL methods. Safe RL methods, without significant modification, cannot directly apply to LTL tasks and instead only address tasks with stationary safety constraints. Offline methods assume access to a fixed dataset and are not designed for the online learning setting we address. Multi-task methods [vaezipoor2021ltl2action, yalcinkaya2023automata, yalcinkaya2024compositional, jackermeier2025deepltl] take special measures to generalize across a distribution of LTL specifications. Comparisons to these methods in a single-task context would therefore require removing significant components from them and would not yield strong conclusions regarding their relative merit. We instead compared our algorithm against two methods for online RL with singular LTL specifications: Counterfactual Experiences for Reward Machines with Reward Shaping (CRM-RS) [toroicarteRewardMachines2022] and the LOF [araki2021logicaloptionsframework]. CRM-RS employs an RM to define the task and dispense shaped rewards that incentivize task completion. RM states in non-accepting sink components are treated as terminal states, implementing a rollout-terminating strategy to enforce safety constraints. The LOF is a framework for learning hierarchical policies to satisfy LTL tasks. It learns a separate policy, referred to as a logical option, for satisfying every subgoal-proposition in the task. When training each logical option, each step that violates the task’s safety propositions receives a large reward penalty (set to $R_{s}=-1000$ based on their implementation) to discourage unsafe behavior. These policies are then composed with a higher-level policy obtained using value iteration over a discrete state space derived from the task automaton and all the subgoal states of the environment. Both these approaches can handle general LTL specifications and, to our knowledge, represent the most relevant baselines to compare our algorithm against. Furthermore, we note that RMs and the LOF cannot be easily enhanced with high-performing GCRL and Safe RL techniques without adopting substantial modifications to their learning algorithms and product MDP formulation. Our method effectively integrates these components in a single framework that enables learning from general logical objectives in complex settings, as demonstrated in the following tasks.
@@
We chose five distinct LTL tasks and three different agents to facilitate a thorough comparison between our algorithm and the two baseline methods. All environment simulation was done within the Brax physics simulator [brax2021github]. The agents used in our experiments are a 2D PointMass, a Quadcopter, and an 8-DOF Ant quadruped. Additional details for these environments are given in the appendix. We used open environments without physical obstacles for our experiments. Instead, obstacles were introduced solely through the task specification. This was done to demonstrate the effectiveness of our algorithm in avoiding obstacles based purely on feedback from the task specifications, without confounding effects such as physically restricted motion. The LTL tasks used in the evaluation include: (1) a two-subgoal sequential navigation task between opposite corners of the environment ($\lozenge(g_{1}\wedge\circ(\lozenge g_{2}))$), (2) a two-subgoal branching navigation task between opposite corners of the environment ($\lozenge g_{1}\wedge\lozenge g_{2}$), (3) a single-goal navigation task constrained by an unsafe region ($\lozenge g_{1}\wedge\square\neg o_{1}$), (4) a two-subgoal navigation task with a disappearing safety constraint ($\neg o_{1}\mathcal{U}g_{1}\wedge\circ\lozenge g_{2}$), and (5) an infinitely-looping navigation task with a persistent safety constraint ($\square\lozenge(g_{1}\wedge\circ\lozenge g_{2})\wedge\square\neg o_{1}$). The starting position, obstacle and subgoal configurations for each environment are shown in Figure LABEL:fig:environments.
@@ Results
The results for our three simulated environments and five task types are shown in Table LABEL:tab:task-experiments. All results are reported for the final policy obtained after 5 million environment interactions with five different seeds. For LOF, each individual logical option was afforded its own 5 million training steps for fairness. For each seed, we evaluated the final policy in 16 randomly initialized episodes lasting 1000 steps. We report the reward, which corresponds to the number of steps spent near the final goal of the task automaton, and the success rate ($S.R.$), which is the proportion of policy rollouts that completely satisfied the task specification for all 1000 steps. Because the robustness of rollouts for the final “Loop” task cannot be meaningfully evaluated over a finite trajectory [hahn2018omegaregularobjectivesmodelfreereinforcement], we instead compute the success rate by examining the proportion of rollouts that are never unsafe and successfully complete at least 1 full loop within the episode.
@@ 5.3 Ablative Analysis
Our second ablation demonstrates the benefit of our minimum safety constraint (2) over the discounted sum-of-cost framework (1) for accurately learning the state-action safety function. We modified our CMDP to provide positive costs when there was a safety violation, trained $Q^{c}_{\theta}$ to predict discounted sums of costs, and constrained $\pi$ to choose actions below an upper limit on total cost $\mathcal{L}$. This limit was chosen based on the best performing value $\mathcal{L}\in\{0,10,40,60\}$. We report the difference in performance over only the tasks that involved safety constraints and observe that our CMDP formulation is critical to the performance of our algorithm. We also show a visualization of the policy differences that result from this ablation in Figure 3. We observe that in our formulation (left), ACQL can more effectively learn the safety constraint, which substantially improves the quality of $Q^{r}_{\psi}$ (bottom). Note that the differing scale for $Q^{c}_{\theta}$ in our formulation (top left) results in a yellow safe region, while the standard formulation depicts low-cost, safe regions in blue.